OpenStudy (anonymous):

If I'm given the coordinates (4,6) and (-1,2) my equation was written out like this... 2-6 m=---- -1-4 I did the equation but got m= -4/-5 What did I do wrong?? I cant figure it out. I DONT want the answer just a brief explanation, please.

OpenStudy (anonymous):

you did everything right you just forgot to cancel your negative for your numerator and denominator so instead of -4/-5 you would have 4/5 :)

OpenStudy (anonymous):

Oh man... Thank you :) Haha, I never know whether I got it right or not (I SUCK at math!!) Haha... Thanks again :)

OpenStudy (anonymous):

no problem :)

OpenStudy (anonymous):

So, if I were to write out the equation would I use the y= mx+b?

OpenStudy (anonymous):

yes you can you can use the slope-intercept form which is y=mx+b or you can use the point-slope form but the slope-intercept form is easier :) =b

OpenStudy (anonymous):

y= mx+b where m= 4/5 and your x,y is one point so you can choose which ever point you want

OpenStudy (anonymous):

okay one more question.. haha... with the given coordinates (above) it would be y=4/5x+ B.... How can I find B? haha

OpenStudy (anonymous):

ohhhhhhh so it can be any y intercept for b?

OpenStudy (anonymous):

no you would solve b by using one of you points and plugging in the x and y value so if we use the point (4,6) we would get y=mx+b where y = 6, m= 4/5 and x = 4 then solve for b 4=4/5(6)+b 4-4/5(6) = b 4-24/5 = b -4/5 = b

OpenStudy (anonymous):

Ohhhhhh I gotcha now! (sorry, my brain is fried right now haha been doing math since 5 this morning!) Haha

OpenStudy (anonymous):

no problem, anytime, glad to help :D

OpenStudy (anonymous):

Good to know! You should have that lifesaver title hahaha

OpenStudy (anonymous):

looool

OpenStudy (anonymous):

haha.... :) well, you totally rock!

OpenStudy (anonymous):

thanks :)

OpenStudy (anonymous):

No problemo :)